# TinyML NLP Approach for Semantic Wireless Sentiment Classification

### v1

@@ IV Endline
We explored TinyML approaches for Semantic text sentiment classification via FL and SL. Our approaches are designed to be energy-efficient and privacy preserving alternatives to CL. It is clear that quantization techniques such as 8-bit quantization appeared to be optimum in our scenario. SL excels in reducing user-side computation and CO2 emmisions, while maintaining accuracy in noisy conditions but incurs higher communication energy. FL, especially with Q8 quantization, offers an optimal balance between computational efficiency, communication cost, and data privacy, making it ideal for decentralized environments with limited bandwidth. The proposed TinyML SL and FL approaches are considered very efficient for tiny devices with limited energy resources, where SL is the most energy efficient and environment friendly. Future work could explore differential privacy to improve efficiency and security.

### v2

@@ IV Conclusion
We explored TinyML approaches for semantic text sentiment classification via FL and SL. Our approaches are designed to be energy-efficient and privacy-preserving alternatives to CL. It is clear that quantization techniques such as 8-bit quantization appeared to be optimum in our scenario. SL excels in reducing user-side computation and CO2 emissions, while maintaining accuracy in noisy conditions but incurs higher communication energy. FL, especially with Q8 quantization, offers an optimal balance between computational efficiency, communication cost, and data privacy, making it ideal for decentralized environments with limited bandwidth. The proposed semantic TinyML SL and FL approaches are considered very efficient for tiny devices with limited energy resources, where SL is the most privacy-preserving, energy-efficient and environment-friendly. Future work could extend these TinyML schemes to LLMs and integrate differential privacy to further enhance communication efficiency and security.
Acknowledgments
This work is supported by the KAUST Office of Sponsored Research under Award ORA-CRG2021-4695.